# Let LLMs Speak Embedding Languages: Generative Text Embeddings via Iterative Contrastive Refinement

### v1

@@ Abstract
Existing large language model (LLM)-based embeddings typically adopt an encoder-only paradigm, treating LLMs as static feature extractors and overlooking their core generative strengths. We introduce GIRCSE (Generative Iterative Refinement for Contrastive Sentence Embeddings), a novel framework that leverages autoregressive generation to iteratively refine semantic representations. By producing sequences of soft tokens optimized under contrastive objective, GIRCSE captures latent concepts and implicit semantics that encoder-only methods often miss. To guide this process, we propose an Iterative Contrastive Refinement (ICR) objective that encourages each refinement step to yield better representations. Extensive experiments show that GIRCSE outperforms strong LLM-based embedding baselines on the MTEB benchmark and instruction-following tasks. Moreover, GIRCSE exhibits an emergent test-time scaling property: generating more tokens at inference steadily improves embedding quality. Our results establish generative iterative refinement as a new paradigm for representation learning.
@@ 1 Introduction
However, current LLM-based embeddings typically adopt an encoder-style design: embeddings are extracted in a single forward pass with contrastive learning objectives, without leveraging the generative capacity of LLMs. This overlooks a key strength of pretrained LLMs—their ability to reason and iteratively refine through autoregressive generation Wei et al. (2022); Muennighoff et al. (2025). This raises a fundamental question: Can LLM-based embedding models also benefit from iterative generation? We hypothesize that generation enables iterative refinement of embeddings, allowing models to progressively consolidate semantics over multiple steps rather than encoding all semantics in a single pass.
@@ 2 Related Work
Early Embedding Models. Text embedding methods have evolved from traditional word-level representations to sophisticated neural approaches. Early methods like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) captured basic semantic relationships but lacked contextual understanding. The introduction of transformer-based models marked a significant advancement, with BERT-based approaches like Sentence-BERT Reimers & Gurevych (2019) and SimCSE Gao et al. (2021) establishing contrastive learning as the dominant paradigm for learning sentence representations.
@@
Table 1: Comparison of LLM-based embedding models. ”Bidir.” indicates bidirectional attention, ”GP” means generated tokens pooling, ”TTS” refers to test-time scaling capability.
@@ 3.1 Generative Embedding Framework
where $p_{\psi}$ represents the generative distribution and $\mathbf{S}_{<k}=(\mathbf{s}_{1},\ldots,\mathbf{s}_{k-1})$ are the previously generated soft tokens. The soft tokens are then mapped into embedding space [^1] We defer the detailed soft token generation mechanism to Section 3.2, while here we focus on the overall framework, producing $\mathbf{D}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{k})\in\mathbb{R}^{K\times d}$, and is concatenated with input embeddings $\mathbf{X}$ to further feed into the Transformer encoder $f_{\theta}$:
@@ 4.1 Experiment Setup
Comparison Methods. We compare GIRCSE against four categories of text embedding models. (1) Non-LLM methods including encoder-based models such as E5-Large Wang et al. (2022), GTE-Large Li et al. (2023b), BGE-Large Li et al. (2023a), and UAE-Large Li & Li (2024). (2) LLM-based methods are instruction-tuned LLM embeddings, including LLM2Vec BehnamGhader et al. (2024), GritLM Muennighoff et al. (2024), E5-Mistral Wang et al. (2024), NV-Embed-v1 Lee et al. (2024), SFR-Embedding-2 Meng* et al. (2024), and gte-Qwen2 Li et al. (2023b). (3) Generative embeddings cover (i) two-stage approaches that expand text with an auxiliary LLM before re-encoding (see Appendix G for detail) and (ii) the end-to-end generative model Inbedder Peng et al. (2024). (4) Fair Baselines are included by re-implementing two paradigms on the same training data for fair comparison: (i) Causal-EOS (causal attention + EOS pooling) and (ii) Bidirectional-Avg (bidirectional attention + average pooling), equivalent to E5-Mistral and GritLM respectively but trained with less data.
@@ 4.2 Main Results
MTEB (English, v2) Instruct Following
Overall
Rank
@@ Appendix C Hyperparameter Settings
Appendix D Pseudo Code for GIRCSE
Algorithm 1 provides a pseudo-code summary of GIRCSE. We include this block to clarify the core computation steps and facilitate reproducibility. The algorithm specifies how input tokens are embedded, how soft embeddings are autoregressively generated across multiple steps, and how the final representation is obtained through pooling. While implementation details may vary, this summary highlights the essential components needed to reproduce our method.
@@ Appendix D Pseudo Code for GIRCSE
Appendix E Scalability Analysis
@@
Appendix F Theoretical Computational and Memory Cost Analysis
@@
Appendix G Detail Implementation of Two-stage Generation Embedding

### v2

@@ Abstract
Existing large language model (LLM)-based embeddings typically adopt an encoder-only paradigm, treating LLMs as static feature extractors and overlooking their core generative strengths. We introduce GIRCSE (Generative Iterative Refinement for Contrastive Sentence Embeddings), a novel framework that leverages autoregressive generation to iteratively refine semantic representations. By producing sequences of soft tokens optimized under contrastive objective, GIRCSE captures latent concepts and implicit semantics that encoder-only methods often miss. To guide this process, we propose an Iterative Contrastive Refinement (ICR) objective that encourages each refinement step to yield better representations. Extensive experiments show that GIRCSE outperforms strong LLM-based embedding baselines on the MTEB benchmark and instruction-following tasks. Moreover, GIRCSE exhibits an emergent test-time scaling property: generating more tokens at inference steadily improves embedding quality. Our results establish generative iterative refinement as a new paradigm for representation learning. Our code and pre-trained models are available at https://github.com/Roytsai27/GIRCSE.
@@ 1 Introduction
However, current LLM-based embeddings typically operate as single-pass feature extractors: embeddings are extracted in a single forward pass with contrastive learning objectives, without leveraging the generative capacity of LLMs. This overlooks a key strength of pretrained LLMs—their ability to reason and iteratively refine through autoregressive generation Wei et al. (2022); Muennighoff et al. (2025). This raises a fundamental question: Can LLM-based embedding models also benefit from iterative generation? We hypothesize that generation enables iterative refinement of embeddings, allowing models to progressively consolidate semantics over multiple steps rather than encoding all semantics in a single pass.
@@ 2 Related Work
Early Embedding Models. Text embedding methods have evolved from traditional word-level representations to sophisticated neural approaches. Early methods like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) captured basic semantic relationships but lacked contextual understanding. The introduction of transformer-based models marked a significant advancement, with BERT-based approaches like Sentence-BERT Reimers and Gurevych (2019) and SimCSE Gao et al. (2021) establishing contrastive learning as the dominant paradigm for learning sentence representations.
@@
Table 1: Comparison of LLM-based embedding models. “Bidir.” indicates bidirectional attention, “GP” means generated tokens pooling, “TTS” refers to test-time scaling capability.
@@ 3.1 Generative Embedding Framework
where $p_{\psi}$ represents the generative distribution and $\mathbf{S}_{<k}=(\mathbf{s}_{1},\ldots,\mathbf{s}_{k-1})$ are the previously generated soft tokens. The soft tokens are then mapped into embedding space [^1] We defer the detailed soft token generation mechanism to Section 3.2, while here we focus on the overall framework, producing $\mathbf{D}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{k})\in\mathbb{R}^{K\times d}$, and is concatenated with input embeddings $\mathbf{X}$ to further feed into the Transformer decoder $f_{\theta}$:
@@ 3.3 Guiding Generative Embedding with Iterative Contrastive Refinement
Output: Final embedding $\mathbf{z}\in\mathbb{R}^{d}$
1ex
$\mathbf{X}\leftarrow(\mathbf{E}[t_{1}],\ldots,\mathbf{E}[t_{N}])$
// Embed the input tokens
$\mathbf{D}\leftarrow[\,]$
// Initialize the list of generated embeddings
for $k=1$ to $K$ do
 $\mathbf{H}\leftarrow f_{\theta}([\mathbf{X};\mathbf{D}])$
 // Forward with input and generated tokens
 $\mathbf{h}^{\prime}_{k-1}\leftarrow\mathbf{H}[N+k-1]$
 // Last hidden state for next-token prediction
 // Compute soft token distribution
 $\mathbf{d}_{k}\leftarrow\sum_{i=1}^{|\mathcal{V}|}s_{k,i}\mathbf{e}_{i}$
 // Compute soft embedding
 $\mathbf{D}\leftarrow\mathbf{D}\|\mathbf{d}_{k}$
 // Append soft embedding for next step
$\mathbf{H}\leftarrow f_{\theta}([\mathbf{X};\mathbf{D}])$
$\mathbf{G}\leftarrow(\mathbf{H}[N+1],\ldots,\mathbf{H}[N+K])$
// Collect the last $K$ hidden states
$\mathbf{z}\leftarrow\mathcal{P}(\mathbf{G})$
// Pool generated representations into a single embedding
return $\mathbf{z}$
Algorithm 1 GIRCSE: Autoregressive Generative Embedding
@@
We summarize the overall training procedure in Algorithm 2. During each training iteration, we sample a mini-batch of queries and documents. For every input sequence, the model performs autoregressive soft token generation for $K$ steps to obtain the sequence of embeddings. Finally, the model parameters $\psi$ are updated end-to-end via gradient descent to minimize the combined objective $\mathcal{L}_{\text{total}}$.
@@ 4.1 Experiment Setup
Comparison Methods. We compare GIRCSE against four categories of text embedding models. (1) Non-LLM methods including encoder-based models such as E5-Large Wang et al. (2022), GTE-Large Li et al. (2023b), BGE-Large Li et al. (2023a), and UAE-Large Li and Li (2024). (2) LLM-based methods are instruction-tuned LLM embeddings, including LLM2Vec BehnamGhader et al. (2024), GritLM Muennighoff et al. (2024), E5-Mistral Wang et al. (2024), NV-Embed-v1 Lee et al. (2024), SFR-Embedding-2 Meng* et al. (2024), and gte-Qwen2 Li et al. (2023b). (3) Generative embeddings cover (i) two-stage approaches that expand text with an auxiliary LLM before re-encoding (see Appendix F for detail) and (ii) the end-to-end generative model Inbedder Peng et al. (2024). (4) Fair Baselines are included by re-implementing two paradigms on the same training data for fair comparison: (i) Causal-EOS (causal attention + EOS pooling) and (ii) Bidirectional-Avg (bidirectional attention + average pooling), equivalent to E5-Mistral and GritLM respectively but trained with less data.
@@ 4.2 Main Results
MTEB (English, v2) Instruct Following Overall Rank
@@ 7 Conclusion
Acknowledgments
This material is based upon work supported by National Science and Technology Council, ROC under grant number 114-2221-E-002-134-MY3 and by Taiwan Centers of Excellence (TCE)
@@ Appendix C Hyperparameter Settings
Appendix D Scalability Analysis
@@
Appendix E Theoretical Computational and Memory Cost Analysis
@@
Appendix F Detail Implementation of Two-stage Generation Embedding
@@
Appendix G Training Stability Analysis
A potential concern when optimizing models that involve the generation of soft tokens is the risk of gradient instability. To empirically validate the stability of our proposed model, GIRCSE, we monitor its training dynamics. We present the training loss curve in Figure 4 and the L2 norm of the gradients in Figure 5.
As shown in the figures, the training loss (Figure 4) demonstrates a smooth and consistent decrease, indicating stable convergence. Furthermore, the gradient norm (Figure 5) converges normally throughout the training process.
Figure 4: Training loss curve of GIRCSE.
Figure 5: Training gradient norm (L2) of GIRCSE, plotted with the top 2% of outliers removed for clarity.